OpenStudy (anonymous):

the length of a rectangular garden is 7 m greater than the width. The area is 78m2. find the length and width

OpenStudy (anonymous):

A=l*w 72=w*(7+w) 72=7w+w^2 w^2+7w-71=0 can you do it from now?

OpenStudy (anonymous):

sorry 78

OpenStudy (anonymous):

hmm. is that to find both?

OpenStudy (anonymous):

A=l*w 78=w*(7+w) 78=7w+w^2 w^2+7w-78=0

OpenStudy (anonymous):

factor will find w

OpenStudy (anonymous):

im really bad at factoring.

OpenStudy (anonymous):

(w+13)(w-6) w=-13 or w=6 the measure can not negative than width =6 , length =6+7=13 check: 13*6 =78

OpenStudy (anonymous):

so length is 13 and w is 78??

OpenStudy (anonymous):

wait, width is 6

OpenStudy (anonymous):

length is 6+7 so its 13

OpenStudy (anonymous):

w for width l for length width =6 length =13 Area = 6*13=78

OpenStudy (anonymous):

length is 6+7 so its 13, yes

OpenStudy (anonymous):

ahhhh yes. thaks